A subway trauma accelerates from rest at one station at a rate of 1.2 m/s^2 for half of the distance to the next station, then decelerates at this same rat for the final half. If the stations are 1100 m apart, find: a. The time of travel between stations b. the maximum speed of the train.
Answer #1

Distance between train stations L=1100\,m

Initially train is at rest.

Now the train accelerates at a=1.2\,m/s^2 till distance covered is \frac{L}{2}=550\,m

using the formula s=ut+\frac{1}{2}at^2

\frac{L}{2}=0+\frac{1}{2}at^2

time taken is t=\sqrt{\frac{L}{a}}=\sqrt{\frac{1100}{1.2}}=30.3\,s

using the formula v^2-u^2=2as

v^2-o^2=2a\times\frac{L}{2}

Final speed achieved is (at half way point) v=\sqrt{aL}=36.3\,m/s

now the train decelerates from this speeed to rest with deceleration a=1.2\,m/s^2 till distance covered is \frac{L}{2}=550\,m

final speed is zero.

using the formula v=u+at

0=\sqrt{aL}-at

time taken is t=\sqrt{\frac{L}{a}}

a. Hence total time of travel between the stations is 60.6 s

b. maximum speed is 36.3 m/s